OpenStudy (anonymous):

Write an equation of a line parallel to line GH, G(-2,2) H(2,6) below in slope intercept form that passes through the point (-5, 6)

OpenStudy (texaschic101):

first we find the slope slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (6 - 2) / (2 - (-2) slope(m) = 4/4 = 1 now we use y = mx + b slope(m) = 1 (-5,6)...x = -5 and y = 6 now we sub and solve for b, the y intercept 6 = 1(-5) + b 6 = -5 + b 6 + 5 = b 11 = b so your slope intercept form is : y = x + 11

OpenStudy (texaschic101):

sorry about before...I read the question wrong

OpenStudy (texaschic101):

questions ?